OpenStudy (anonymous):

through (2,-5) ; perpendicular to 3y=x-6 I got y=3x-11 the answer is y=-3x-1. what did i do wrong

OpenStudy (anonymous):

okay, so perpendicular to 3y = x- 6 means that the slope of the desired equation is the negative reciprocal of 3y = x-6. However this equation is not yet in proper form. To get it in the form y = mx + b, we divide the equation by 3 on both sides, to get y = (x/3) - 2. From this we get that the slope is 1/3. So the slope of the equation you are deriving is -3, because -3 * 1/3 = -1. So, y = -3x + b. To find b, the y intercept, plug in the (x, y) pair (2, -5) you were told the equation goes through. -5 = -3*2 + b; b = -5 + 6 = 1; so y = -3x + 1

OpenStudy (anonymous):

you went wrong when you said the y-intercept was -1, because it is in fact 1

OpenStudy (anonymous):

y=(1/3) x-2 perpendicular to slope is negative reciprocal so u did 3 where it is supposed to be -3

OpenStudy (anonymous):

uh ok goddess explained it

OpenStudy (anonymous):

where did -3 come from

OpenStudy (anonymous):

From the answer you derived it seems you know the general method, you just have to be careful with the arithmatic

OpenStudy (anonymous):

if m is slope of a line then its peripendicular is -1/m

OpenStudy (anonymous):

-3 is the negative reciprocal of 1/3. The slope of an equation perpendicular to another equation is always the negative reciprocal of the slope of the given equation

OpenStudy (anonymous):

so if m=1/2 perpendicular is -2

OpenStudy (anonymous):

to find the negative reciprocal, set your given * a variable = -1. For example, x*1/3 = -1; x = -1/(1/3) = -1 * 3/1 = -3

OpenStudy (anonymous):

Hope this helps

OpenStudy (anonymous):

its confusing when ur teacher doesnt even explain this in the first place

OpenStudy (anonymous):

It is unfortunate that some teachers don't explain things well. Try going back to your textbook or check online for explanations. There are often very clear ones available.

OpenStudy (anonymous):

i learned all of physics on wikipedia

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i think teachers explain how but not usually the why and that screws every1 over when they are attempting problems